# A Survey on Knowledge Distillation of Large Language Models

### v3

@@ 3.5K Wikipedia Categories +
RLCD Yang et al. (2024a) IF/Preference Human-written Prompts LLaMA LLaMA Labeling SFT + RL
@@ 4.2.2 Preference
The previously mentioned methods primarily focus on the basic capability of student models to produce outcomes that are strictly accurate but may not align with human preferences, reaching alignment at this level enables these models to aid in various tasks without meeting higher-level demands. Early methods mainly utilize human feedback for the alignment of human preferences Ziegler et al. (2019); Stiennon et al. (2020); Wu et al. (2021); Ouyang et al. (2022); Bai et al. (2022b); Köpf et al. (2023); Yuan et al. (2023b). However, obtaining human feedback is costly and labor-intensive, thus methods that learn from AI feedback are also proposed to align with human preferences Bai et al. (2022a); Kwon et al. (2023); Scheurer et al. (2023); Kim et al. (2023a); Roit et al. (2023); Yang et al. (2024a); Lee et al. (2023a); Tunstall et al. (2023); Cui et al. (2023a); Wang et al. (2023f).
The concept of RLAIF, introduced by Bai et al. (2022a), involves the integration of preferences labeled by LLMs with those labeled by humans. This approach is designed to simultaneously optimize two key objectives: ensuring the helpfulness of the output and minimizing any potential harm, making the responses of LLMs more aligned with Human preferences. Kwon et al. (2023) develop a proxy reward function using LLMs like GPT-3, which is created by first providing the LLM with a description of the behaviors desired by the user, along with a small number of examples. The LLM then produces rewards by evaluating how closely the outputs of a model align with the provided descriptions, essentially measuring their relevance to the established ground truth. Scheurer et al. (2023) propose Imitation Learning from Language Feedback, in which a language model is utilized to improve various outputs generated by a model. This refinement is based on a reference provided by a human. Following this process, the most effectively refined output is chosen to be used in further supervised fine-tuning. As outlined by Kim et al. (2023a), ALMoST involves condensing human preferences into a set of heuristic guidelines. An example of such a rule is the idea that larger LLMs that utilize more comprehensive and higher-quality prompts are likely to yield superior responses. Based on these established guidelines, comparison data is generated using responses from LLMs of different sizes and with varying prompts. This data is then used to train a reward model. Yang et al. (2024a) propose Reinforcement Learning from Contrast Distillation, which aims to align language models without relying on human feedback. This approach involves training a preference model using simulated pairs of preferences, including both high-quality and low-quality examples which are generated through contrasting prompts, positive and negative.
@@ 4.2.3 Value
Though slightly under-explored, aligning LLMs with human values by distilling is still possible Bai et al. (2022a); Cui et al. (2023a); Yang et al. (2024a); Sun et al. (2024b). For instance, Bai et al. (2022a) propose RLAIF, utilizing AI-generated labels to interactively improve both helpfulness and harmlessness. Sun et al. (2024b) prompt the student model with 16 principles as guidelines for generating helpful, ethical, and reliable responses. Similarly, both harmless and harmful generations could be elicited by modifying the prompts, and then are used to train the preference model Yang et al. (2024a). Cui et al. (2023a) utilize GPT-4 to rank generations regarding helpfulness, truthfulness, and honesty. Liu et al. (2023b) advance the alignment of LLMs with societal values by incorporating simulated social interactions into the training process. This approach encompasses a range of elements, including demonstrations that are both in alignment and in conflict with social norms, as well as collective ratings, in-depth feedback, and responses that are revised iteratively.
@@ 4.4.5 Text Generation Evaluation
Text generation evaluation, i.e. NLG evaluation, focuses on assessing the quality of generated content. Unlike traditional NLG evaluation metrics like BLEU Papineni et al. (2002) or ROUGE Lin (2004), which primarily rely on surface-level text comparisons, LLMs, trained on extensive corpora and refined through techniques like RLHF, offer a more nuanced and human-aligned assessment. This sophistication has led to the increasing use of LLMs in NLG evaluation (detailed further in  Li et al. (2024b)). Through KD of LLMs, student evaluators could enhance inference efficiency and achieve more flexible and highly customized evaluation Wang et al. (2023b); Kim et al. (2024); Xu et al. (2023d); Jiang et al. (2023c); Li et al. (2024a).
PandaLM Wang et al. (2023b) concentrates on a pairwise evaluator designed to compare two pieces of generated content. It utilizes a teacher LLM (GPT-3.5) to judge which response is better for a given instruction and input, providing reasons for its decision. Addressing the need for customized and flexible criteria to meet realistic user demands, Prometheus Kim et al. (2024) distills GPT-4 to construct a training dataset that includes reference answers and a variety of customized scoring rubrics. This dataset is then used to tune LLaMA for evaluating model-generated responses. Instructscore Xu et al. (2023d) takes a more fine-grained approach by using GPT-4 to create detailed analysis data. This data is employed to tune LLaMA, enabling it to perform error analysis on generated texts compared to reference texts. The system further refines its evaluation capabilities through self-training with real model-generated response-reference pairs. For reference-free evaluation across diverse domains, TigerScore Jiang et al. (2023c) samples data from a variety of text generation datasets, such as summarization, translation, and data-to-text. It distills error analysis knowledge from GPT-4 and uses this to fine-tune LLaMA for more nuanced evaluation. Lastly, to adapt evaluation to real-world scenarios beyond conventional NLP tasks, Auto-J Li et al. (2024a) collects real-world user queries and their evaluations from a teacher LLM. This massive dataset of real-world scenarios is then used to distill evaluation knowledge into LLaMA through fine-tuning, enhancing its practical applicability.
@@ 5.1 Law
Law holds a crucial position in molding societies, overseeing human interactions, and ensuring justice prevails. Informed decision-making, legal interpretation, and the provision of legal advice by professionals hinge on precise and current information. Legal intelligent applications in different scenarios usually require combinations of multiple fundamental capabilities of legal text retrieval, understanding, reasoning and generating Zhang et al. (2023g); Sun (2023); Lai et al. (2023). Addressing the intricacies of legal terminology, subtle interpretations, and the constant evolution of legislation presents distinctive challenges that demand customized resolutions. To handle the above challenges, several studies have investigated the customization of LLMs for intelligent legal services Cui et al. (2023b); Yue et al. (2023b); Huang et al. (2023b); Wu et al. (2023d). This involves a continued pre-training process on extensive legal corpora, followed by fine-tuning with self-constructed instructions or augmented data using advanced LLMs.
@@ 5.2 Medical and Healthcare
The integration of LLMs carries substantial promise in fundamentally reshaping the landscape of medical data analysis, comprehension, and smart medical services (Singhal et al., 2023; Yang et al., 2024b). Significant research endeavors have been dedicated to adapting general-purpose LLMs to the medical domain, given the ever-expanding wealth of information encompassing electronic health records, medical literature, and clinical data. Especially in healthcare, LLMs are revolutionizing patient care, research, and administrative efficiency. They enhance diagnostic accuracy by analyzing patient data and medical literature, offering personalized recommendations, and identifying potential drug interactions. LLMs also streamline administrative tasks by automating patient documentation and processing insurance claims, reducing the burden on healthcare providers and improving patient experiences. Furthermore, they facilitate medical research by synthesizing vast amounts of data to uncover new insights into diseases and treatments (will be discussed later). This integration of LLMs into healthcare is paving the way for more informed clinical decision-making, improved patient outcomes, and more efficient healthcare systems.
These adaptations extend across a spectrum, ranging from refining the precision of medical diagnoses Wang et al. (2023k) and providing personalized treatment recommendations Zhu et al. (2023) to automating routine administrative processes within healthcare settings.
@@
While existing studies predominantly concentrate on training using dedicated medical dialogue datasets comprising medical textbooks Wu et al. (2023e), biomedical papers Luo et al. (2023e) medical knowledge-graphs Bao et al. (2023b), or authentic doctor-patient interactions Bao et al. (2023b), an expanding body of research is delving into the augmentation of medical instruction-following data with advanced LLMs to enhance the alignment of the intricacies within practical user instructions. Zhang et al. (2023c) introduce HuatuoGPT specifically tailored for medical consultations. The model leverages both distilled data from ChatGPT and real-world data from doctors during the supervised fine-tuning stage. In a parallel effort, Xiong et al. (2023) construct a dataset of medical dialogues in Chinese, employing ChatGPT’s assistance. Their methodology encompassed various techniques to train DoctorGLM, an easily deployable LLM designed for tasks such as diagnoses, drug recommendations, and other medical advice. Zhang et al. (2023f) fine-tune LLaMA-series models using 52k diverse, machine-generated, medical instruction-following data named MedInstruct-52k. This effort resulted in the development of AlpaCare, a model demonstrating robust medical proficiency and generalizability across both general and medical-specific domain free-form instruction evaluations. In a different vein, Wang et al. (2023a) propose HuaTuo, a LLaMA-based model that undergoes supervised fine-tuning with generated QA instances. This refinement process enhances the model’s possession of more reliable medical knowledge. Li et al. (2023i) introduce ChatDoctor, which was first trained as a generic conversation model based on LLaMA. It utilized 52K instruction-following data from Stanford University’s Alpaca project Taori et al. (2023). Subsequently, the conversation model underwent fine-tuning on a dataset of 100K patient-physician conversations collected from an online medical consultation website. This two-step training process underscores the model’s adaptability to diverse conversational contexts, particularly those specific to patient-physician interactions.
@@ 5.3 Finance
The application of LLMs to the finance domain (Xue et al., 2023) significantly transforms how financial data is analyzed, decisions are made, and customer interactions are managed. In finance, LLMs offer unprecedented capabilities in understanding complex financial documents, predicting market trends, and automating risk assessment, thus enabling more informed and faster decision-making processes. By processing and analyzing vast amounts of unstructured financial data, such as news articles, reports, and real-time market feeds, LLMs can identify patterns and insights that were previously inaccessible, leading to more accurate forecasts and strategic financial planning. Furthermore, LLMs enhance customer experiences through personalized financial advice, automated customer service, and sophisticated chatbots that can handle intricate queries. This level of automation and insight has the potential to increase efficiency, reduce operational costs, and improve compliance and risk management practices in financial institutions, making LLMs a transformative force in the finance sector.
Knowledge distillation from a proprietary LLM is still under-explored, and most existing works focus on adapting LLMs to finance applications by continual pre-training on finance-specific corpora (Wu et al., 2023g; Lu et al., 2023) or fine-tuning in a supervised manner on multi-task finance-specific instructions (Yang et al., 2023e; Xie et al., 2023b; Wang et al., 2023l).
@@ 5.4 Science
The integration of LLMs into the science domain (Taylor et al., 2022; Yin et al., 2023b) represents a paradigm shift in research, knowledge discovery, and the dissemination of scientific information. In science, LLMs are leveraged to digest and synthesize vast amounts of literature, aiding in the identification of new research opportunities and the acceleration of scientific breakthroughs. They facilitate the understanding of complex scientific concepts by summarizing research papers, generating hypotheses, and even drafting research proposals and manuscripts, thus significantly reducing the time researchers spend on literature review and enabling them to focus more on experimental work. LLMs also democratize access to scientific knowledge by providing layperson summaries of complex research findings, making science more accessible to non-experts and fostering a broader public understanding of scientific advancements. By enhancing the efficiency of research workflows and fostering interdisciplinary collaborations, LLMs are poised to accelerate the pace of scientific discovery and innovation across various fields. To distill knowledge from an LLM, DARWIN Series (Xie et al., 2023a) utilizes a semi self-instruct for instruction generation for science papers, which is then used to fine-tune an LLM. SciGLM (Zhang et al., 2024) proposes to train a scientific LLM, which prompts a teacher LLM to generate detailed answers for unlabelled scientific questions, as well as a self-reflective critic-and-revise to improve data quality.
Besides the above knowledge distillation methods to adapt LLMs to science, we will also delve into how the distillation happens in sub-domains, e.g., mathematics, astronautics, chemistry, etc.
@@
Biology.
In the field of Biology, particularly in the study of proteins, DNA, and RNA, LLMs are revolutionizing our understanding of the fundamental molecules of life. By analyzing vast datasets of biological sequences and structures, LLMs can predict the three-dimensional shapes of proteins, potential functions, and interactions at a scale and speed beyond traditional computational methods. This capability is critical for unraveling the complexities of biological systems, advancing drug discovery by identifying targets and designing molecules with high precision, and understanding genetic diseases through the interpretation of genomic variations.
@@
Geography, Geology, and Environmental Science. The integration of LLMs into Geography, Geology, and Environmental Science is revolutionizing these fields by enhancing data analysis, predictive modeling, and interdisciplinary research Roberts et al. (2023); Lin et al. (2023b); Wang et al. (2023m).
K2 (Deng et al., 2023), the first-ever LLM specialized in the geoscience domain, demonstrates the significant impact of knowledge distillation in vertical domain specialization. By adapting the general-domain LLaMA-7B model with a 5.5B token geoscience corpus and introducing the GeoSignal instruction tuning dataset, K2 showcases enhanced performance in geoscience knowledge understanding and utilization. The model’s development highlights a novel approach to efficiently gather domain-specific data and align model responses to specialized user queries, underpinning the importance of domain-specified vertical distillation in advancing research and applications within geoscience. OceanGPT (Bi et al., 2023), introduced as the first LLM for ocean science tasks, underscores the vital role of knowledge distillation in the vertical domain of oceanography. It leverages DOINSTRUCT, a novel framework for generating domain-specific instruction data through multi-agent collaboration, and establishes OCEANBENCH, a benchmark for evaluating LLMs in the ocean domain. The model’s comprehensive experiments demonstrate its superior capability in understanding and generating knowledge for ocean science, showcasing the significant potential of targeted knowledge distillation in enhancing domain-specific model performance. MarineGPT (Zheng et al., 2023b) showcases the transformative potential of knowledge distillation in the marine domain by leveraging a novel vision-language model tailored for marine science. Utilizing the Marine-5M dataset, which includes over 5 million marine image-text pairs, MarineGPT excels in providing detailed, accurate, and domain-specific responses. This advancement underscores the model’s ability to significantly enhance marine knowledge comprehension and application, emphasizing the critical role of domain-specific distillation in bridging the gap between general-purpose models and specialized domain requirements. GeoGalactica (Lin et al., 2024) represents a pioneering step in specializing LLMs for geoscience, leveraging a 30 billion parameter model pre-trained on a vast geoscience corpus. This model, notable for being the largest of its kind within the geoscience domain, showcases the significant potential of knowledge distillation in fostering scientific discoveries by bridging artificial intelligence with geoscience research and applications.

### v4

@@ 3.5K Wikipedia Categories +
RLCD Yang et al. (2024) IF/Preference Human-written Prompts LLaMA LLaMA Labeling SFT + RL
@@ 4.2.2 Preference
The previously mentioned methods primarily focus on the basic capability of student models to produce outcomes that are strictly accurate but may not align with human preferences, reaching alignment at this level enables these models to aid in various tasks without meeting higher-level demands. Early methods mainly utilize human feedback for the alignment of human preferences Ziegler et al. (2019); Stiennon et al. (2020); Wu et al. (2021); Ouyang et al. (2022); Bai et al. (2022b); Köpf et al. (2023); Yuan et al. (2023b). However, obtaining human feedback is costly and labor-intensive, thus methods that learn from AI feedback are also proposed to align with human preferences Bai et al. (2022a); Kwon et al. (2023); Scheurer et al. (2023); Kim et al. (2023a); Roit et al. (2023); Yang et al. (2024); Lee et al. (2023a); Tunstall et al. (2023); Cui et al. (2023a); Wang et al. (2023f).
The concept of RLAIF, introduced by Bai et al. (2022a), involves the integration of preferences labeled by LLMs with those labeled by humans. This approach is designed to simultaneously optimize two key objectives: ensuring the helpfulness of the output and minimizing any potential harm, making the responses of LLMs more aligned with Human preferences. Kwon et al. (2023) develop a proxy reward function using LLMs like GPT-3, which is created by first providing the LLM with a description of the behaviors desired by the user, along with a small number of examples. The LLM then produces rewards by evaluating how closely the outputs of a model align with the provided descriptions, essentially measuring their relevance to the established ground truth. Scheurer et al. (2023) propose Imitation Learning from Language Feedback, in which a language model is utilized to improve various outputs generated by a model. This refinement is based on a reference provided by a human. Following this process, the most effectively refined output is chosen to be used in further supervised fine-tuning. As outlined by Kim et al. (2023a), ALMoST involves condensing human preferences into a set of heuristic guidelines. An example of such a rule is the idea that larger LLMs that utilize more comprehensive and higher-quality prompts are likely to yield superior responses. Based on these established guidelines, comparison data is generated using responses from LLMs of different sizes and with varying prompts. This data is then used to train a reward model. Yang et al. (2024) propose Reinforcement Learning from Contrast Distillation, which aims to align language models without relying on human feedback. This approach involves training a preference model using simulated pairs of preferences, including both high-quality and low-quality examples which are generated through contrasting prompts, positive and negative.
@@ 4.2.3 Value
Though slightly under-explored, aligning LLMs with human values by distilling is still possible Bai et al. (2022a); Cui et al. (2023a); Yang et al. (2024); Sun et al. (2024b). For instance, Bai et al. (2022a) propose RLAIF, utilizing AI-generated labels to interactively improve both helpfulness and harmlessness. Sun et al. (2024b) prompt the student model with 16 principles as guidelines for generating helpful, ethical, and reliable responses. Similarly, both harmless and harmful generations could be elicited by modifying the prompts, and then are used to train the preference model Yang et al. (2024). Cui et al. (2023a) utilize GPT-4 to rank generations regarding helpfulness, truthfulness, and honesty. Liu et al. (2023b) advance the alignment of LLMs with societal values by incorporating simulated social interactions into the training process. This approach encompasses a range of elements, including demonstrations that are both in alignment and in conflict with social norms, as well as collective ratings, in-depth feedback, and responses that are revised iteratively.
@@ 4.4.5 Text Generation Evaluation
Text generation evaluation, i.e. NLG evaluation, focuses on assessing the quality of generated content. Unlike traditional NLG evaluation metrics like BLEU Papineni et al. (2002) or ROUGE Lin (2004), which primarily rely on surface-level text comparisons, LLMs, trained on extensive corpora and refined through techniques like RLHF, offer a more human-aligned assessment. This sophistication has led to the increasing use of LLMs in NLG evaluation (detailed further in  Li et al. (2024b)). Through KD of LLMs, student evaluators could enhance inference efficiency and achieve more flexible and highly customized evaluation Wang et al. (2023b); Kim et al. (2024); Xu et al. (2023d); Jiang et al. (2023c); Li et al. (2024a).
PandaLM Wang et al. (2023b) concentrates on a pairwise evaluator designed to compare two pieces of generated content. It utilizes a teacher LLM (GPT-3.5) to judge which response is better for a given instruction and input, providing reasons for its decision. Addressing the need for customized and flexible criteria to meet realistic user demands, Prometheus Kim et al. (2024) distills GPT-4 to construct a training dataset that includes reference answers and a variety of customized scoring rubrics. This dataset is then used to tune LLaMA for evaluating model-generated responses. Instructscore Xu et al. (2023d) takes a more fine-grained approach by using GPT-4 to create detailed analysis data. This data is employed to tune LLaMA, enabling it to perform error analysis on generated texts compared to reference texts. The system further refines its evaluation capabilities through self-training with real model-generated response-reference pairs. For reference-free evaluation across diverse domains, TigerScore Jiang et al. (2023c) samples data from a variety of text generation datasets, such as summarization, translation, and data-to-text. It distills error analysis knowledge from GPT-4 and uses this to fine-tune LLaMA. Lastly, to adapt evaluation to real-world scenarios beyond conventional NLP tasks, Auto-J Li et al. (2024a) collects real-world user queries and their evaluations from a teacher LLM. This massive dataset of real-world scenarios is then used to distill evaluation knowledge into LLaMA through fine-tuning, enhancing its practical applicability.
@@ 5.1 Law
Law holds a crucial position in molding societies, overseeing human interactions, and ensuring justice prevails. Informed decision-making, legal interpretation, and the provision of legal advice by professionals hinge on precise and current information. Legal intelligent applications in different scenarios usually require combinations of multiple fundamental capabilities of legal text retrieval, understanding, reasoning and generating Zhang et al. (2023g); Sun (2023); Lai et al. (2023). To address challenges like legal terminology, subtle interpretations, and the constant evolution of legislation presents distinctive challenges that demand customized resolutions. To handle the above challenges, several studies have investigated the customization of LLMs for intelligent legal services Cui et al. (2023b); Yue et al. (2023b); Huang et al. (2023b); Wu et al. (2023d). This involves a continued pre-training process on extensive legal corpora, followed by fine-tuning with self-constructed instructions or augmented data using advanced LLMs.
@@ 5.2 Medical and Healthcare
The integration of LLMs holds great potential for transforming medicine and healthcare. Extensive research has focused on adapting general-purpose LLMs to the medical domain Singhal et al. (2023), such as electronic health records, and healthcare applications like patient care Zhu et al. (2023). Recent work has focused on enhancing medical instruction-following data with advanced teacher LLMs to better align with complex user instructions. Given the abundance of medical data, most studies combine real-world data with distilled instruction data from teacher LLMs Zhang et al. (2023c); Xiong et al. (2023); Zhang et al. (2023f); Wang et al. (2023a); Li et al. (2023i); Han et al. (2023); Wu et al. (2023f); Bao et al. (2023a); Chen et al. (2023d).
@@
While existing studies predominantly concentrate on training using dedicated medical dialogue datasets comprising medical textbooks Wu et al. (2023e), biomedical papers Luo et al. (2023e) medical knowledge-graphs Bao et al. (2023b), or authentic doctor-patient interactions Bao et al. (2023b), an expanding body of research is delving into the augmentation of medical instruction-following data with advanced LLMs to enhance the alignment with practical user instructions. Zhang et al. (2023c) introduce HuatuoGPT specifically tailored for medical consultations. The model leverages both distilled data from ChatGPT and real-world data from doctors during the supervised fine-tuning stage. In a parallel effort, Xiong et al. (2023) construct a dataset of medical dialogues in Chinese, employing ChatGPT’s assistance. Their methodology encompassed various techniques to train DoctorGLM, an easily deployable LLM designed for tasks such as diagnoses, drug recommendations, and other medical advice. Zhang et al. (2023f) fine-tune LLaMA-series models using 52k diverse, machine-generated, medical instruction-following data named MedInstruct-52k. This effort resulted in the development of AlpaCare, a model demonstrating robust medical proficiency and generalizability across both general and medical-specific domain free-form instruction evaluations. In a different vein, Wang et al. (2023a) propose HuaTuo, a LLaMA-based model that undergoes supervised fine-tuning with generated QA instances. This refinement process enhances the model’s possession of more reliable medical knowledge. Li et al. (2023i) introduce ChatDoctor, which was first trained as a generic conversation model based on LLaMA. It utilized 52K instruction-following data from Stanford University’s Alpaca project Taori et al. (2023). Subsequently, the conversation model underwent fine-tuning on a dataset of 100K patient-physician conversations collected from an online medical consultation website. This two-step training process underscores the model’s adaptability to diverse conversational contexts, particularly those specific to patient-physician interactions.
@@ 5.3 Finance
The application of LLMs to the finance domain (Xue et al., 2023) significantly transforms how financial data is analyzed, decisions are made, and customer interactions are managed. In finance, LLMs offer unprecedented capabilities in understanding complex financial documents, predicting market trends, and automating risk assessment, thus enabling more informed and faster decision-making processes. By processing and analyzing vast amounts of unstructured financial data, such as news articles, reports, and real-time market feeds, LLMs can identify patterns and insights that were previously inaccessible, leading to more accurate forecasts and strategic financial planning. Furthermore, LLMs enhance customer experiences through personalized financial advice, automated customer service, and sophisticated chatbots that can handle complex queries. This level of automation and insight has the potential to increase efficiency, reduce operational costs, and improve compliance and risk management practices in financial institutions, making LLMs a transformative force in the finance sector. Knowledge distillation from a proprietary LLM is still under-explored, and most existing works focus on adapting LLMs to finance applications by continual pre-training on finance-specific corpora (Wu et al., 2023g; Lu et al., 2023) or fine-tuning in a supervised manner on multi-task finance-specific instructions (Yang et al., 2023e; Xie et al., 2023b; Wang et al., 2023k).
@@ 5.4 Science
The integration of LLMs into the science domain (Taylor et al., 2022; Yin et al., 2023b) represents a paradigm shift in research, knowledge discovery, and the dissemination of scientific information. In science, LLMs are leveraged to digest and synthesize vast amounts of literature, aiding in the identification of new research opportunities and the acceleration of scientific breakthroughs. They facilitate the understanding of complex scientific concepts by summarizing research papers, generating hypotheses, and even drafting research proposals and manuscripts, thus significantly reducing the time researchers spend on literature review and enabling them to focus more on experimental work. LLMs also democratize access to scientific knowledge by providing layperson summaries of complex research findings, making science more accessible to non-experts and fostering a broader public understanding of scientific advancements. By enhancing the efficiency of research workflows and fostering interdisciplinary collaborations, LLMs are poised to accelerate the pace of scientific discovery and innovation across various fields. To distill knowledge from an LLM, DARWIN Series (Xie et al., 2023a) utilizes a semi self-instruct for instruction generation for science papers, which is then used to fine-tune an LLM. SciGLM (Zhang et al., 2024) proposes to train a scientific LLM, which prompts a teacher LLM to generate detailed answers for unlabelled scientific questions, as well as a self-reflective critic-and-revise to improve data quality. Besides the above knowledge distillation methods to adapt LLMs to science, we will also delve into how the distillation happens in sub-domains, e.g., mathematics, astronautics, chemistry, etc.
@@
Biology. In the field of Biology, particularly in the study of proteins, DNA, and RNA, LLMs are revolutionizing our understanding of the fundamental molecules of life. By analyzing vast datasets of biological sequences and structures, LLMs can predict the three-dimensional shapes of proteins, potential functions, and interactions at a scale and speed beyond traditional computational methods. This capability is critical for unraveling the complexities of biological systems, advancing drug discovery by identifying targets and designing molecules with high precision, and understanding genetic diseases through the interpretation of genomic variations.
@@
Geography, Geology, and Environmental Science. The integration of LLMs into Geography, Geology, and Environmental Science is revolutionizing these fields by enhancing data analysis, predictive modeling, and interdisciplinary research Roberts et al. (2023); Lin et al. (2023b); Wang et al. (2023l).
K2 (Deng et al., 2023), the first-ever LLM specialized in the geoscience domain, demonstrates the significant impact of knowledge distillation in vertical domain specialization. By adapting the general-domain LLaMA-7B model with a 5.5B token geoscience corpus and introducing the GeoSignal instruction tuning dataset, K2 showcases enhanced performance in geoscience knowledge understanding and utilization. The model’s development highlights a novel approach to efficiently gather domain-specific data and align model responses to specialized user queries. OceanGPT (Bi et al., 2023), introduced as the first LLM for ocean science tasks, underscores the vital role of knowledge distillation in the vertical domain of oceanography. It leverages DOINSTRUCT, a novel framework for generating domain-specific instruction data through multi-agent collaboration, and establishes OCEANBENCH, a benchmark for evaluating LLMs in the ocean domain. MarineGPT (Zheng et al., 2023b) showcases the transformative potential of knowledge distillation in the marine domain by leveraging a novel vision-language model tailored for marine science. Utilizing the Marine-5M dataset, which includes over 5 million marine image-text pairs, MarineGPT excels in providing detailed, accurate, and domain-specific responses. GeoGalactica (Lin et al., 2024) represents a pioneering step in specializing LLMs for geoscience, leveraging a 30 billion parameter model pre-trained on a vast geoscience corpus. This model is notable for being the largest of its kind within the geoscience domain.